OpenStudy (anonymous):

The area of a rectangle is the product of width and length. If a rectangle has length 15m sqrt 9m and width sqrt 3m , which of the following is the area of the rectangle?

OpenStudy (anonymous):

A = W * L = √3 * 15 √9 = 45 √3

OpenStudy (anonymous):

thank you, i was able to check my answer but i couldnt figure out if it was 45m^2 \[\sqrt{3}\] or 45m\[\sqrt{3}\]m

OpenStudy (anonymous):



OpenStudy (anonymous):

= or 45m 3√ m